OpenStudy (gamenerd123):

A recursive rule for a geometric sequence is a1=4/9; an=3(an-1) What is an explicit rule for this sequence?

OpenStudy (gamenerd123):

An =?

OpenStudy (3mar):

Are you familiar with that?

OpenStudy (gamenerd123):

Yes

OpenStudy (gamenerd123):

But when it comes to recursive terms, its a little confusing for me

OpenStudy (3mar):

No worries.. Step by step. \[\Huge a_1=\frac{ 4 }{ 9 }~~~,~~~ a_n=3*a_{n-1}\] That means two things: - the first term in this sequence is 4/9 - the common ration is 3, i.e to get any term you have to multiply the previous term by 3 to get the new term agree so far?

OpenStudy (gamenerd123):

Yes, I catch on so far

OpenStudy (gamenerd123):

What do I do next?

OpenStudy (3mar):

That is great! Now for the explicit rule for this sequence: you need to type out the first term \(\color{red}{a_1}\) in the sequence followed by the common ration \(\color{red}r\) raised to the power \(\color{red}{n-1}\) Why???? because when you are going to find out the first term, you would plug in n=1 into this \(a_n=a_1*r^{n-1}\), so you got the first term \(a_1\) only, so you can try yourself any other term

OpenStudy (gamenerd123):

So what will that mean?

OpenStudy (gamenerd123):

For the problem

OpenStudy (3mar):

\[\Huge a_n=\color{blue}{a_1}*\color{lime}r^{n-1}\] that means you plug in (4/9) instead of "the first term" \(a_1\) and 3 instead of the common ratio \(r\)

OpenStudy (gamenerd123):

So how will that turn out?

OpenStudy (3mar):

It is your turn @Gamenerd123 now

OpenStudy (gamenerd123):

Well, i mean like how do I do that, sorry.

OpenStudy (3mar):

Where is the first term? and where is the common ratio?

OpenStudy (gamenerd123):

a1 is the first term and 3 is the common ratio, right?

OpenStudy (3mar):

a1=??

OpenStudy (gamenerd123):

3?

OpenStudy (3mar):

if a1=3 and r=3, so 4/9=?????????????????

OpenStudy (gamenerd123):

3

OpenStudy (3mar):

a_1=? r=?

OpenStudy (gamenerd123):

Its not 3?

OpenStudy (3mar):

no it is not.. \[\Huge a_1=\frac{ 4 }{ 9 }~~~,~~~ a_n=3*a_{n-1}\] Do you remember that?

OpenStudy (gamenerd123):

Yes

OpenStudy (3mar):

so?

OpenStudy (gamenerd123):

Hmmm...

OpenStudy (3mar):

I really like who ask and stop the explanation when he sticks to something, even it looks so simple! So don't be embarrassed if there is anything not clear for you

OpenStudy (gamenerd123):

My minds so slow for math I'm sorry

OpenStudy (gamenerd123):

It takes a while for me to get it

OpenStudy (3mar):

No no don't say that ...I am not agree with you!

OpenStudy (3mar):

I have another idea...

OpenStudy (gamenerd123):

Please, do tell..

OpenStudy (3mar):

Ok.. Can you type the formula of the explicit rule for this sequence?

OpenStudy (gamenerd123):

Not exactly...

OpenStudy (gamenerd123):

I'm very dumb for formulas

OpenStudy (3mar):

 got it?